# A Black Hole Airy Tail

### v1

@@ Abstract
In Jackiw-Teitelboim (JT) gravity, which is dual to a random matrix ensemble, the annealed entropy differs from the quenched entropy at low temperatures and goes negative. However, computing the quenched entropy in JT gravity requires a replica limit that is poorly understood. To circumvent this, we define an intermediate quantity called the semi-quenched entropy, which has the positivity properties of the quenched entropy, while requiring a much simpler replica trick. We compute this in JT gravity in different regimes using i) a bulk calculation involving wormholes corresponding to the Airy limit of the dual matrix integral and ii) a boundary calculation involving one-eigenvalue instantons, demonstrating consistency between these two calculations in their common regime of validity. We also clarify why a recent attempt to compute the quenched entropy using one-eigenvalue instantons is unreliable due to a breakdown of the saddle-point approximation for the one-eigenvalue instanton in the replica limit.
@@ I Introduction
However, calculating the quenched entropy is not easy. To compute the quenched entropy, one uses the so-called no-replica trick: $\langle\log Z(\beta)\rangle=\partial_{m}\langle Z(\beta)^{m}\rangle\evaluated{}_{m=0}$. At integer $m>0$, this is an $m$-boundary calculation which can be done using the gravitational path integral and involves wormholes connecting different boundaries. This calculation was attempted [3, 10], but the analytic continuation in $m$ is complicated, and so far, there is no general method to compute $S_{Q}$ in gravitational theories. Recently, [4] also attempted a calculation of $S_{Q}$ using the RMT dual to JT gravity. The proposal was that a new saddle, a one-eigenvalue instanton, dominates in the RMT due to the insertion of the operator $\langle Z(\beta)^{m}\rangle$ in Eq. (2) for large $\beta$ [^2] This new saddle in the matrix integral can potentially be described by a bulk saddle involving a semiclassical brane dual to the one-eigenvalue instanton. Such a description is enticing since the properties of the black hole microstate ground state (in a sector of fixed charge) can be captured through a semiclassical description [4]. However, we will point out issues with the replica limit in this calculation that make it unsatisfactory.
@@
Our second goal is to clarify the behavior of the quenched entropy in this low-temperature regime. Critically analyzing the proposal of [4], we show in Sec. IV that while the one-eigenvalue instanton provides the correct answer for $\langle Z(\beta)^{m}\rangle$ for $m\neq 0$, the analytic continuation to $m=0$ required for the quenched entropy invalidates the result: in the limit $m\to 0$, the one-eigenvalue instanton reunites with the rest of the spectrum and the saddle-point approximation in its derivation breaks down due to a noncommuting order of limits between $S_{0}\to\infty$ and $m\to 0$. We demonstrate that the one-eigenvalue instanton answer in fact disagrees with previous matrix model calculations of the quenched entropy in the low temperature limit [11, 12]. Consequently, the one-eigenvalue instanton provides the correct result for computationally simpler quantities like the annealed and semi-quenched entropy, but not for the quenched entropy.
@@ III.1 Airy Wormholes at $\beta=O(e^{2S_{0}/3})$
fixed. In this limit, $Z\left(\beta\right)$ is dominated by eigenvalues near the edge of the spectrum. From the perspective of RMT, it is well known that the square-root edge of the spectrum, Eq. (3), is controlled by the Airy model [13], see [4] for a recent review.
@@
The $l=0$ term dominates in our limit of interest. Remarkably, for $l=0$, the sum over genus is convergent and can be explicitly performed; see [3] for a review. The geometries dominating this sum are well understood [14], and the resulting partition function matches the Airy matrix model [15]  [^4] This limit has also been used recently in understanding the spectral form factor in [16, 17]. See also [18] from a different bulk interpretation of the genus resummation..
@@
where $\text{disk}_{\text{res}}\left(x\right)$ is the one-boundary partition function with boundary length $x$ and $\text{cyl}_{\text{res}}\left(x,y\right)$ is the connected two-boundary partition function with boundary lengths $x,y$. In this limit, both $\text{disk}_{\text{res}}\left(x\right)$ and $\text{cyl}_{\text{res}}\left(x,y\right)$ receive important contributions from higher-genus geometries. These quantities can be computed exactly by resumming the genus expansion and the relevant formulas are [15, 3]:
@@ III.2 Eigenvalue Instantons at $\beta\gg O(e^{2S_{0}/3})$
We discuss aspects of the general solution in Appendix B. To compare the instanton answer with the Airy answer, we focus on the regime $\beta=\alpha e^{2S_{0}/3}$ with $\alpha=O(1)$ but large. Since the remaining eigenvalues remain in their original saddle, we will focus on the effective action for the lowest eigenvalue given by  [^6] Note that [4] uses a different normalization compared to the usual JT gravity density of states. The answers are related by rescaling both $\beta$ and $e^{S_{0}}$ by a factor of 2.
@@ IV Comments on Quenched Entropy
At $\beta\gtrsim e^{2S_{0}/3}$, the answer for the quenched entropy is also controlled by the Airy limit [11, 12]. Due to the replica limits involved in this calculation, it is much harder to obtain directly but a matrix model analysis was done in [11] where they found that the low temperature entropy behaves like $T^{1+\upbeta}$, where $\upbeta=1,2,4$ is the Dyson index that characterizes the GOE, GUE and GSE ensembles  [^8] The $T^{1+\upbeta}$ behavior of the quenched entropy holds at arbitrarily low temperatures, up to doubly non-perturbative corrections [11].. In particular, the leading order term at large $\alpha$ is $S_{Q}^{(n)}\approx\frac{7\pi^{4}\gamma^{3}}{360\alpha^{3}}\frac{(1+n)(1+n^{2})}{n^{3}}$. The quenched entropy, obtained numerically by sampling 1000 GUE $1000\times 1000$ matrices, is plotted in Fig. 1 and matches the aforementioned analytic result at large $\alpha$.
In contrast, the proposal of [4] using the one-eigenvalue instanton fails to reproduce this, leading to $S_{Q}\approx T^{3/2}$ at low temperatures, independent of the symmetry class. Here, we explain why this analysis fails.
As seen above, the location of the instanton is a function of $m$ and, unlike the semi-quenched entropy, approaches the continuum edge as we take $m\to 0$. In this limit, the saddle point approximation breaks down since higher loop contributions in Eq. (22) become important. One should have always been suspicious of this analysis since the leading answer comes from the continuum spectrum in this limit, whereas the assumption used in obtaining the instanton was that the leading eigenvalue dominates $Z\left(\beta\right)$. This is a consequence of taking the wrong order of limits between $\beta\to\infty$ and $m\to 0$  [^9] This is similar in spirit to issues arising for Rényi entropies for states where fluctuations in the geometry are large [20, 21]..
Comparing to Sec. III, the key difference for the semi-quenched and annealed entropies is that they are dominated at large $\beta$ by configurations with large gaps between the two leading eigenvalues, whereas the quenched entropy gets its leading (power-law) contribution from configurations with a small gap.
@@ Acknowledgements.
We thank Sergio Hernandez-Cuenca, Geoff Penington, and Wayne Weng for helpful discussions. This work was supported in part by the Leinweber Institute for Theoretical Physics, by the Department of Energy, Office of Science, Office of High Energy Physics through DE-SC0025522, DE-SC0019380, DE-SC0025293, and DE-FOA-0002563, by AFOSR award FA9550-22-1-0098, and by a Sloan Fellowship. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Semi-quenched Rényi entropy at higher $n$
$\begin{split}&\mathcal{E}^{(m)}(\alpha_{1},\ldots,\alpha_{m})\equiv\frac{\exp\left(\sum_{i=1}^{m}\alpha_{i}^{3}/24\right)}{(4\pi)^{m/2}\sqrt{\prod_{i=1}^{m}x_{i}}}\\ \times&\int_{s_{i}\geq 0}d^{m}s\,\exp\left(-\frac{1}{2}\sum_{i=1}^{m}\left(\frac{(s_{i}-s_{i+1})^{2}}{\alpha_{i}}-(s_{i}+s_{i+1})\alpha_{i}\right)\right)\end{split}$ (24)

### v3

@@ Abstract
In Jackiw-Teitelboim (JT) gravity, which is dual to a random matrix ensemble, the annealed entropy differs from the quenched entropy at low temperatures and goes negative. However, computing the quenched entropy in JT gravity requires a replica limit that is poorly understood. To circumvent this, we define an intermediate quantity called the semi-quenched entropy, which has the positivity properties of the quenched entropy, while requiring a much simpler replica trick. We compute this in JT gravity in different regimes using i) a bulk calculation involving wormholes corresponding to the Airy limit of the dual matrix integral and ii) a boundary calculation involving one-eigenvalue instanton saddles proposed by Hernández-Cuenca, demonstrating consistency between these two calculations in their common regime of validity. We also clarify why similar one-eigenvalue instanton saddles cannot be used to compute the quenched entropy due to a breakdown of the saddle-point approximation for the one-eigenvalue instanton in the replica limit.
@@ I Introduction
However, calculating the quenched entropy is not easy. To compute the quenched entropy, one uses the so-called no-replica trick: $\langle\log Z(\beta)\rangle=\partial_{m}\langle Z(\beta)^{m}\rangle\evaluated{}_{m=0}$. At integer $m>0$, this is an $m$-boundary calculation which can be done using the gravitational path integral and involves wormholes connecting different boundaries. This calculation was attempted [3, 10], but the analytic continuation in $m$ is complicated, and so far, there is no general method to compute $S_{Q}$ in gravitational theories. Recently, [4], among other interesting results, proposed a calculation of $S_{Q}$ using the RMT dual to JT gravity. The interesting proposal was that a new saddle, a one-eigenvalue instanton, dominates in the RMT due to the insertion of the operator $\langle Z(\beta)^{m}\rangle$ in Eq. (2) for large $\beta$ [^2] This new saddle in the matrix integral can potentially be described by a bulk saddle involving a semiclassical brane dual to the one-eigenvalue instanton. Such a description is enticing since the properties of the black hole microstate ground state (in a sector of fixed charge) can be captured through a semiclassical description [4]. The one-eigenvalue instanton indeed correctly computes $\langle Z(\beta)^{m}\rangle$ and, consequently, the annealed entropy. However, we will point out issues with the replica limit when calculating the quenched entropy using this technique.
@@
Our second goal is to clarify the behavior of the quenched entropy in this low-temperature regime. In Sec. IV we show that while the one-eigenvalue instanton provides the correct answer for $\langle Z(\beta)^{m}\rangle$ for $m\neq 0$, the analytic continuation to $m=0$ required for the quenched entropy is subtle: in the limit $m\to 0$, the one-eigenvalue instanton reunites with the rest of the spectrum and the saddle-point approximation in its derivation breaks down due to a noncommuting order of limits between $S_{0}\to\infty$ and $m\to 0$. We demonstrate that the one-eigenvalue instanton answer disagrees with previous matrix model calculations of the quenched entropy in the low temperature limit [11, 12]. Thus, the one-eigenvalue instanton provides the correct result for computationally simpler quantities like the annealed and semi-quenched entropy, but not for the quenched entropy.
@@ III.1 Airy Wormholes at $\beta=O(e^{2S_{0}/3})$
fixed. In this limit, $Z\left(\beta\right)$ is dominated by eigenvalues near the edge of the spectrum. From the perspective of RMT, it is well known that the square-root edge of the spectrum, Eq. (3), is controlled by the Airy model [13], see [4] for a nice review.
@@
The $l=0$ term dominates in our limit of interest. Remarkably, for $l=0$, the sum over genus is convergent and can be explicitly performed; see [3, 14] for a review. The geometries dominating this sum are well understood [15], and the resulting partition function matches the Airy matrix model [16]  [^4] This limit has also been used recently in understanding the spectral form factor in [17, 18]. See also [14] for a different bulk interpretation of the genus resummation..
@@
where $\text{disk}_{\text{res}}\left(x\right)$ is the one-boundary partition function with boundary length $x$ and $\text{cyl}_{\text{res}}\left(x,y\right)$ is the connected two-boundary partition function with boundary lengths $x,y$. In this limit, both $\text{disk}_{\text{res}}\left(x\right)$ and $\text{cyl}_{\text{res}}\left(x,y\right)$ receive important contributions from higher-genus geometries. These quantities can be computed exactly by resumming the genus expansion and the relevant formulas are [16, 19, 20, 3]:
@@ III.2 Eigenvalue Instantons at $\beta\gg O(e^{2S_{0}/3})$
We discuss aspects of the general solution in Appendix B. To compare the instanton answer with the Airy answer, we focus on the regime $\beta=\alpha e^{2S_{0}/3}$ with $\alpha=O(1)$ but large. Since the remaining eigenvalues remain in their original saddle, we will focus on the effective action for the lowest eigenvalue given by
@@ IV Comments on Quenched Entropy
At $\beta\gtrsim e^{2S_{0}/3}$, the answer for the quenched entropy is also controlled by the Airy limit [11, 12]. Due to the replica limits involved in this calculation, it is much harder to obtain directly but a matrix model analysis was done in [11] where they found that the low temperature entropy behaves like $T^{1+\upbeta}$, where $\upbeta=1,2,4$ is the Dyson index that characterizes the GOE, GUE and GSE ensembles  [^7] This behavior holds at arbitrarily low temperatures, up to doubly non-perturbative corrections [11].. In particular, for GUE, the leading order term at large $\alpha$ is $S_{Q}^{(n)}\approx\frac{7\pi^{4}\gamma^{3}}{360\alpha^{3}}\frac{(1+n)(1+n^{2})}{n^{3}}$. The quenched entropy, obtained numerically by sampling 1000 GUE $\times 1000$ 1000 matrices, is plotted in Fig. 1 and matches the aforementioned analytic result at large $\alpha$.
In contrast, the proposal of using the one-eigenvalue instanton for the quenched entropy [4] does not reproduce this, leading to $S_{Q}\approx T^{3/2}$ at low temperatures, independent of the symmetry class.
As seen above, the location of the instanton is a function of $m$ and, unlike the semi-quenched entropy, approaches the continuum edge as we take $m\to 0$. In this limit, the saddle point approximation breaks down since higher-loop contributions in Eq. (22) become important. One should have always been suspicious of this analysis since the leading answer comes from the continuum spectrum in this limit, whereas the assumption used in obtaining the instanton was that the leading eigenvalue dominates $Z\left(\beta\right)$. This is a consequence of taking the wrong order of limits between $\beta\to\infty$ and $m\to 0$  [^8] This is similar in spirit to issues arising for Rényi entropies for states where fluctuations in the geometry are large [22, 23]..
Compared to Sec. III, the key difference for the semi-quenched and annealed entropies is that they are dominated at large $\beta$ by configurations with large gaps between the two leading eigenvalues, whereas the quenched entropy gets its leading (power-law) contribution from configurations with a small gap.
@@ Acknowledgements.
We are especially grateful to Sergio Hernández-Cuenca for numerous discussions. We also thank Geoff Penington and Wayne Weng for helpful discussions. This work was supported in part by the Leinweber Institute for Theoretical Physics, by the Department of Energy, Office of Science, Office of High Energy Physics through DE-SC0025522, DE-SC0019380, DE-SC0025293, and DE-FOA-0002563, by AFOSR award FA9550-22-1-0098, and by a Sloan Fellowship. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Semi-quenched Rényi entropy at higher $n$
$\begin{split}&\mathcal{E}^{(m)}(\alpha_{1},\ldots,\alpha_{m})\equiv\frac{\exp\left(\sum_{i=1}^{m}\alpha_{i}^{3}/24\right)}{(4\pi)^{m/2}\sqrt{\prod_{i=1}^{m}\alpha_{i}}}\\ \times&\int_{s_{i}\geq 0}d^{m}s\,\exp\left(-\frac{1}{2}\sum_{i=1}^{m}\left(\frac{(s_{i}-s_{i+1})^{2}}{\alpha_{i}}-(s_{i}+s_{i+1})\alpha_{i}\right)\right)\end{split}$ (24)